OpenStudy (anonymous):

The equation of line EF is y = 1 over 2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2)

OpenStudy (bibby):

is t hat \(\frac{1}{2x+6}\) or \(\frac{1}{2}x+6\)

OpenStudy (anonymous):

no :( ill put up the answer choices

OpenStudy (anonymous):

y = −2x − 2 y = negative 1 over 2x + 2 y = 1 over 2x − 2 y = − 2x + 2

OpenStudy (bibby):

I'm asking you what the line EF is

OpenStudy (bibby):

the equation*

OpenStudy (anonymous):

the first response you gave me

OpenStudy (bibby):

so \(\frac{1}{2x+6}\)

OpenStudy (anonymous):

yes

OpenStudy (bibby):

wolfram says the slope would be 1, I'll test that out later now we use the point slope form \(y-y_1=m(x-x_1)\) where \((x_1,y_1)\) is the point we want the lines to pass through

OpenStudy (anonymous):

okay so what do I plug in?

OpenStudy (bibby):

m is the slope

OpenStudy (bibby):

and the point we want it to go through is (0, −2)

OpenStudy (bibby):

(0,-2)

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

y-(-2)=1/2(x-0) that's what i have so far plugged in

OpenStudy (bibby):

what I got was \(y-(-2)=1(x-0)\) y+2=x y=x-2

OpenStudy (anonymous):

Thank you! omg I get how you solved it

OpenStudy (bibby):

now I'll do the testing thing to make sure the slope is actually 1

OpenStudy (anonymous):

Um another quick question is 4/-3 a positive slope or a negative slope?

OpenStudy (bibby):

negative. positive/negative and negative/positive are negative

OpenStudy (bibby):

the thing I got isn't even one of the choices, hmm

OpenStudy (anonymous):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2). well how about a question that's similar to the one were trying to figure out?

OpenStudy (bibby):

that one's easier

OpenStudy (bibby):

what is the slope of LM?

OpenStudy (anonymous):